OpenStudy (faiqraees):

Find the remainder when 5^1998 is divided by 11 @ganeshie8

ganeshie8 (ganeshie8):

Let's look for a pattern in the powers 5^1 = `5` 5^2 = 25 = `3` 5^3 = 5^2*5 = 3*5 = 15 = `4` 5^4 = 3^2 = `9` 5^5 = 5^4*5 = 9*5 = 45 = `1` 5^6 = 5^5*5 = 1*5 = `5`

ganeshie8 (ganeshie8):

See the pattern in the remainders of 5^n ? 5, 3, 4, 9, 1, 5, ...

OpenStudy (faiqraees):

So the pattern is 5,3,4,9,1,5 .... and I just have to figure out what the number will be when n is 1998, right thanks

OpenStudy (faiqraees):

Okay one more thing Find the value of m such that the equations x³+mx-1=0 x³-3x+m = 0 have a common root

ganeshie8 (ganeshie8):

Let "a" be the common root.

ganeshie8 (ganeshie8):

Then we have a^3 + ma - 1 = 0 a^3 - 3a + m = 0

ganeshie8 (ganeshie8):

subtract the equations and you can express "m" in terms of "a"

OpenStudy (faiqraees):

x = A [1]-[2] Am-1 +3A -m = 0 A(3+m) = m+1 A = m+1/m+3 3[1] +m[2] 3A³+3Am-3+A³m -3Am+m²=0 A³(3+m)+m²-3=0 A³= (3-m²)/3+m A³= A³ (m+1/m+3)³ = (3-m²)/3+m But this doesnt give me m = -2

OpenStudy (faiqraees):

@ganeshie8

OpenStudy (johnweldon1993):

Are you sure? Try plugging -2 in there again

OpenStudy (faiqraees):

@johnweldon1993 It also gives me some irrational numbers as roots too. What should I do about them?